OpenStudy (anonymous):

Can someone help me with this problem? I thought that you were just supposed to subtract "a-4" from the original problem to result in, 3a-12 but that is incorrect it says. Subtract "a-4" from "4a+16" and simplify your answer. The variable "a" should appear only once in your answer.

OpenStudy (anonymous):

Problem with signs. Your idea is correct what happens when you subtract -4 from 16. \[16-\left( -4 \right) = ?\]

OpenStudy (anonymous):

I guess

OpenStudy (anonymous):

but how should I solve the problem

OpenStudy (anonymous):

All that I know is that 3a+12 is incorrect

OpenStudy (anonymous):

\[\left( 4a+16 \right) - \left( a-4 \right) = 4a+16-a+4 = ?\]What do you get?

OpenStudy (anonymous):

3a+20?

OpenStudy (anonymous):

Correct. Good job. Do you understand where you went wrong the first time?

OpenStudy (anonymous):

Yes I did not correctly follow the second step, the conversion between (a-4) and a+4 thankyou

OpenStudy (anonymous):

You're welcome.